OpenStudy (anonymous):

A baseball team wins 45% of its games. What fraction of games does it win?

OpenStudy (anonymous):

The fraction of games won is 9/20.

OpenStudy (anonymous):

\[\frac{ 45 }{ 100 } or \frac{ 9 }{ 20 }\]

OpenStudy (anonymous):

Please click the medal nextto my answer.

OpenStudy (jiteshmeghwal9):

45%=\(\dfrac{45}{100}\) cancel it :)

OpenStudy (anonymous):

Thanks! :D